"This morning, you borrowed $150,000 to buy a house. The mortgage rate is 7.35 percent. The loan is to be repaid in equal monthly payments over 20 years. The first payment is due one month from today. How much is the equal monthly payment? (Assume that each month is equal to 1/12 of a year." $1,166.52 $1,033.46 $1,102.51 $925.83 $1,194.67 please show all work via a financial calculator preferably TI BA II plus
Answer #1

put the following values in the financial calculator:

N = 20x12 = 240; I/Y = 7.35/12 = 0.6125; PV = 150,000; FV = 0;

Press CPT, then PMT, which gives -1,194.67

Hence, Option "E" is correct.